# Verina: Benchmarking Verifiable Code Generation

### v1

@@ Abstract
Large language models (LLMs) are increasingly integrated in software development, but ensuring correctness in LLM-generated code remains challenging and often requires costly manual review. Verifiable code generation—jointly generating code, specifications, and proofs of code-specification alignment—offers a promising path to address this limitation and further unleash LLMs’ benefits in coding. Yet, there exists a significant gap in evaluation: current benchmarks often lack support for end-to-end verifiable code generation. In this paper, we introduce Verina (Verifiable Code Generation Arena), a high-quality benchmark enabling a comprehensive and modular evaluation of code, specification, and proof generation as well as their compositions. Verina consists of 189 manually curated coding tasks in Lean, with detailed problem descriptions, reference implementations, formal specifications, and extensive test suites. Our extensive evaluation of state-of-the-art LLMs reveals significant challenges in verifiable code generation, especially in proof generation, underscoring the need for improving LLM-based theorem provers in verification domains. The best model, OpenAI o4-mini, generates only 61.4% correct code, 51.0% sound and complete specifications, and 3.6% successful proofs, with one trial per task. We hope Verina will catalyze progress in verifiable code generation by providing a rigorous and comprehensive benchmark. We release our dataset on https://huggingface.co/datasets/sunblaze-ucb/verina and our evaluation code on https://github.com/sunblaze-ucb/verina.
@@ 1 Introduction
Large language models (LLMs) have shown strong performance in programming [1, 2, 3] and are widely adopted in tools like Cursor and GitHub Copilot to boost developer productivity [4]. LLM-generated code is becoming prevalent in commercial software [5] and may eventually form a substantial portion of the world’s code. However, due to their probabilistic nature, LLMs alone cannot provide formal guarantees for the generated code. As a result, the generated code often contains bugs, such as functional errors [6] and security vulnerabilities [7]. When LLM-based code generation is increasingly adopted, these issues can become a productivity bottleneck, as they typically require human review to be resolved [8]. Formal verification presents a promising path to establish correctness guarantees in LLM-generated code but has traditionally been limited to safety-critical applications due to high cost [9, 10, 11]. Similarly to how they scale up code generation, LLMs have the potential to significantly lower the barrier of formal verification. By jointly generating code, formal specifications, and formal proofs of alignment between code and specifications, LLMs can offer higher levels of correctness assurance and automation in software development. This approach represents an emerging programming paradigm known as verifiable code generation [12, 13].
@@
Given the transformative potential of verifiable code generation, it is crucial to develop suitable benchmarks to track progress and guide future development. This is challenging because verifiable code generation involves three interconnected tasks: code, specification, and proof generation. We need to curate high-quality samples and establish robust evaluation metrics for each individual task, while also composing individual tasks to reflect real-world end-to-end usage scenarios where LLMs automate the creation of verified software directly from high-level requirements. Existing benchmarks, as listed in Table 1 and detailed in Section 2, fall short as they lack comprehensive support for all three tasks [14, 15, 16], quality control [17], robust metrics [18], or a modular design [12].
@@
To bridge this gap, we introduce Verina (Verifiable Code Generation Arena), a high-quality benchmark to comprehensively evaluate verifiable code generation. It consists of 189 programming challenges with detailed problem descriptions, code, specifications, proofs, and comprehensive test suites. We format these problems in Lean [19], a general-purpose programming language with a rapidly growing ecosystem and applications in both formal mathematics [20, 21] and verification [22, 23].
@@
Verina is constructed with careful quality control. It draws problems from various sources, including MBPP [18, 24], LiveCodeBench [1], and LeetCode, offering a diverse range of difficulty levels. All samples in the benchmark are manually inspected and revised to ensure clear natural language descriptions and accurate formal specifications and code implementations. Moreover, each sample also includes a comprehensive test suite with both positive and negative cases, which achieves 100% line coverage on the code implementation and passes the ground truth specification.
@@
Verina facilitates the evaluation of code, specification, and proof generation, along with flexible combinations of these individual tasks. We utilize the standard pass@$k$ metric [25] with our comprehensive test suites to evaluate code generation. For proof generation, we use the Lean compiler to automatically verify their correctness. Furthermore, we develop a practical, testing-based approach based on to automatically evaluate model-generated specifications, by verifying their soundness and completeness with respect to ground truth specifications.
@@
The high-quality samples and robust metrics of Verina establish it as a rigorous platform for evaluating verifiable code generation. We conduct a thorough experimental evaluation of nine state-of-the-art LLMs on Verina. Our results reveal that even the top-performing LLM, OpenAI o4-mini, struggles with verifiable code generation, producing only 61.4% correct code solutions, 51.0% sound and complete specifications, and 3.6% successful proof, given a single sample for each task. Interestingly, iterative refinement using Lean compiler feedback can increase the proof success rate to 22.2% with 64 refinement steps. However, this approach significantly raises costs and the 22.2% success rate is still insufficient. These findings underscore the challenges of verifiable code generation and highlight the critical role of Verina in advancing the field.
@@ 2 Background and Related Work
Task support for verifiable code generation. Writing code, specifications, and proofs for a verified software component is time-consuming when done manually. Although various studies have explored using LLMs to automate these tasks, they primarily focus on individual aspects, which are insufficient for fully automating end-to-end verifiable code generation. Benchmarks like HumanEval [3] and MBPP [24] have sparked impressive progress on LLM-based code generation but do not handle formal specifications or proofs. Many verification-focused efforts target only one or two tasks, while assuming the other elements are provided by the human user. For example, DafnyBench [14] and miniCodeProps [26] are two benchmarks designed exclusively for proof generation. Moreover, AutoSpec [27] and SpecGen [28] infer specifications and proofs from human-written code.
@@
To the best of our knowledge, Dafny-Synthesis [18] and Clover [12] are the only two works that cover all three tasks, like Verina. However, they target automated theorem proving using Dafny [29], while Verina leverages interactive theorem proving in Lean. Moreover, they have relatively small numbers of human-written samples (50 and 62 respectively). In contrast, Verina provides 189 high-quality samples with varying difficulty levels.
@@
Automated and interactive theorem proving. A major challenge in formal verification and verifiable code generation lies in tooling. Verification-oriented languages like Dafny [29] and Verus [30] leverage SMT solvers for automated theorem proving [31, 32] and consume only proof hints, such as loop invariants [33] and assertions [34]. However, SMT solvers handle only limited proof domains and behave as black boxes, which can make proofs brittle and hard to debug [35]. Interactive theorem proving (ITP) systems like Lean provide a promising target for verifiable code generation with LLMs. ITPs support constructing proofs with explicit intermediate steps. This visibility enables LLMs to diagnose errors, learn from unsuccessful steps, and iteratively refine their proofs. While ITPs traditionally require humans to construct proofs, recent work shows that LLMs can generate proofs at human level in math competitions [36]. To our knowledge, the only existing verification benchmarks in Lean are miniCodeProps [26] and FVAPPS [17]. miniCodeProps translates 201 Haskell programs and their specifications into Lean but is designed for proof generation only. FVAPPS contains 4,715 Lean programs with LLM-generated specifications from a fully automated pipeline that lacks human validation and quality control. In contrast, Verina provides high-quality, human-verified samples and captures all three foundational tasks in verifiable code generation.
@@
Task compositionality. A key strength of Verina is its modular design, which enables flexible evaluation of not only individual tasks but also their combinations. This compositionality captures diverse real-world scenarios—from specification-guided code generation to end-to-end verifiable code generation—enabling a comprehensive assessment of different aspects of verifiable code generation. This modularity also facilitates targeted research on specific weaknesses, such as improving proof generation. On the contrary, all other prior works lack compositionality. For example, Dafny-Synthesis [18] and Clover [12] mix specification and proof generation into a single task, lacking support for separate evaluation of each.
Table 1: A comparison of Verina with related works on LLMs for code generation and verification. We characterize whether each work supports the three foundational tasks for end-to-end verifiable code generation: CodeGen, SpecGen, ProofGen (Section 4.1). ● means fully supported, ◐ means partially supported, ○ means unsupported. If ProofGen is supported, we specify the proving style: automated theorem proving (ATP) or interactive theorem proving (ITP). For works supporting multiple tasks, we annotate if these tasks are supported in a modular and composable manner. Overall, Verina offers more comprehensive and high-quality benchmarking compared to prior works.
CodeGen SpecGen ProofGen Proving Style Compositionality Language
Benchmarks HumanEval [3], MBPP [24] Python
Dafny-Synthesis [18] ATP Dafny
DafnyBench [14] ATP Dafny
miniCodeProps [26] ITP Lean
FVAPPS [17] ITP Lean
Techniques nl2postcond [37] Python, Java
Clover [12] ATP Dafny
AlphaVerus [15] ATP Rust
AutoSpec [27] ATP C/C++
SpecGen [28] ATP Java
SAFE [16] ATP Rust
AutoVerus [38] ATP Rust
Laurel [34] ATP Dafny
Pei et al. [33] ATP Java
Baldur [39], Selene [40] ITP Isabelle
Rango [41], PALM [42] ITP Coq
Verina ITP Lean
@@
5
6-- Code implementation
7def removeElement (s : Array Int) (k : Nat) (h_precond : removeElement_pre s k) : Array Int :=
@@
9
10-- Pre-condition
11def removeElement_pre (s : Array Int) (k : Nat) : Prop :=
@@
13
14-- Post-condition
15def removeElement_post (s : Array Int) (k : Nat) (result: Array Int) (h_precond : removeElement_pre s k) : Prop :=
16 result.size = s.size - 1 -- Only one element is removed
17 (∀ i, i < k result[i]! = s[i]!) -- The elements before index k remain unchanged
18 (∀ i, i < result.size i k result[i]! = s[i + 1]!) -- The elements after index k are shifted by one position
19
20-- Proof
21theorem removeElement_spec (s: Array Int) (k: Nat) (h_precond : removeElement_pre s k) :
22 removeElement_post s k (removeElement s k h_precond) h_precond := by sorry -- The proof is omitted for brevity
23
24-- Test cases
25(s : #[1, 2, 3, 4, 5]) (k : 2) (result : #[1, 2, 4, 5]) -- Positive test with valid inputs and output
26(s : #[1, 2, 3, 4, 5]) (k : 5) -- Negative test: inputs violate the pre-condition at Line 12
27(s : #[1, 2, 3, 4, 5]) (k : 2) (result : #[1, 2, 4]) -- Negative test: output violates the post-condition at Line 16
28(s : #[1, 2, 3, 4, 5]) (k : 2) (result : #[2, 2, 4, 5]) -- Negative test: output violates the post-condition at Line 17
29(s : #[1, 2, 3, 4, 5]) (k : 2) (result : #[1, 2, 4, 4]) -- Negative test: output violates the post-condition at Line 18
@@ 3.1 Overview and Data Format
  • Code (Line 6–8): ground truth code implementation that solves the programming problem.
  • Specification (Line 10–18): ground truth formal specification for the programming problem. It consists of a pre-condition, which states properties the inputs must satisfy, and a post-condition, which states desired relationship between inputs and outputs.
  • Proof (Optional, Line 20–22): formal proof establishing that the code satisfies the specification. Ground truth proofs are optional in Verina, as they are not required for evaluation. Model-generated proofs can be checked by Lean directly. Nevertheless, we invest significant manual effort in writing proofs for 46 out of 189 examples as they help quality assurance (Section 3.2).
  • Test suite (Line 24–29): a comprehensive suite of both positive and negative test cases. Positive tests are valid input-output pairs that meet both the pre-condition and the post-condition. Negative tests are invalid inputs-output pairs, which means either the inputs violate the pre-condition or the output violates the post-condition. These test cases are useful for evaluating model-generated code and specifications, as detailed in Section 4.1. They are formatted in Lean during evaluation.

### v2

@@ Abstract
Large language models (LLMs) are increasingly integrated in software development, but ensuring correctness in LLM-generated code remains challenging and often requires costly manual review. Verifiable code generation—jointly generating code, specifications, and proofs of code-specification alignment—offers a promising path to address this limitation and further unleash LLMs’ benefits in coding. Yet, there exists a significant gap in evaluation: current benchmarks often focus on only individual components rather than providing a holistic evaluation framework of all tasks. In this paper, we introduce Verina (Verifiable Code Generation Arena), a high-quality benchmark enabling a comprehensive and modular evaluation of code, specification, and proof generation as well as their compositions. Verina consists of 189 manually curated coding tasks in Lean, with detailed problem descriptions, reference implementations, formal specifications, and extensive test suites. Our extensive evaluation of state-of-the-art LLMs reveals significant challenges in verifiable code generation, especially in proof generation, underscoring the need for improving LLM-based theorem provers in verification domains. The best model, OpenAI o4-mini, achieves a 61.4% code correctness rate, 51.0% for specification soundness and completeness, and a mere 3.6% proof success rate (based on one trial per task). We hope Verina will catalyze progress in verifiable code generation by providing a rigorous and comprehensive benchmark. We release our dataset on https://huggingface.co/datasets/sunblaze-ucb/verina and our evaluation code on https://github.com/sunblaze-ucb/verina.
@@ 1 Introduction
Large language models (LLMs) have shown strong performance in programming (Jain et al., 2025; Jimenez et al., 2024; Chen et al., 2021) and are widely adopted in tools like Cursor and GitHub Copilot to boost developer productivity (Kalliamvakou, ). LLM-generated code is becoming prevalent in commercial software (Peters, 2024) and may eventually form a substantial portion of the world’s code. However, due to their probabilistic nature, LLMs alone cannot provide formal guarantees for the generated code. As a result, the generated code often contains bugs, such as functional errors (Wang et al., 2025) and security vulnerabilities (Pearce et al., 2022). When LLM-based code generation is increasingly adopted, these issues can become a productivity bottleneck, as they typically require human review to be resolved (Finley, ). Formal verification presents a promising path to establish correctness guarantees in LLM-generated code but has traditionally been limited to safety-critical applications due to high cost (Gu et al., 2016; Leroy et al., 2016; Bhargavan et al., 2013). Similarly to how they scale up code generation, LLMs have the potential to significantly lower the barrier of formal verification. By jointly generating code, formal specifications, and formal proofs of alignment between code and specifications, LLMs can offer higher levels of correctness assurance and automation in software development. This approach represents an emerging programming paradigm known as verifiable code generation (Sun et al., 2024; Yang et al., 2024).
@@
Given the transformative potential of verifiable code generation, it is crucial to develop suitable benchmarks to track progress and guide future development. This is challenging because verifiable code generation involves three interconnected tasks: code, specification, and proof generation. We need to curate high-quality samples and establish robust evaluation metrics for each individual task, while also composing individual tasks to reflect real-world end-to-end usage scenarios where LLMs automate the creation of verified software directly from high-level requirements. Existing benchmarks, as discussed in Section 2, fall short as they lack comprehensive support for all three tasks (Loughridge et al., 2025; Aggarwal et al., 2024; Chen et al., 2024), quality control (Dougherty & Mehta, 2025), robust metrics (Misu et al., 2024), or a modular design (Sun et al., 2024).
@@
To bridge this gap, we introduce Verina (Verifiable Code Generation Arena), a high-quality benchmark to comprehensively evaluate verifiable code generation. It consists of 189 programming challenges with detailed problem descriptions, code, specifications, proofs, and comprehensive test suites. We format these problems in Lean (Moura & Ullrich, 2021), a general-purpose programming language with a rapidly growing ecosystem and applications in both formal mathematics (Mathlib community, 2020; Mathlib Community, 2022) and verification (de Medeiros et al., 2025; Hietala & Torlak, 2024). Lean has become the one of the most popular platform for LLM-assisted theorem-proving and verification, demonstrated by breakthrough results like AlphaProof (Google DeepMind, 2024) and production adoption at organizations like AWS (de Moura, ), with ongoing efforts to verify mainstream languages like Rust (Ho & Protzenko, 2022).
@@
Verina is constructed with careful quality control. It draws problems from various sources, including MBPP (Misu et al., 2024; Austin et al., 2021), LiveCodeBench (Jain et al., 2025), and LeetCode, offering a diverse range of difficulty levels. All samples in the benchmark are manually inspected and revised to ensure clear text descriptions and accurate formal specifications and code implementations. Moreover, each sample also includes a comprehensive test suite with both positive and negative cases, which achieves 100% code coverage and passes the ground truth specification.
@@
Verina facilitates the evaluation of code, specification, and proof generation, along with flexible combinations of these individual tasks. We utilize the standard pass@$k$ metric (Fan et al., 2024) with our comprehensive test suites to evaluate code generation. For proof generation, we use the Lean compiler to automatically verify their correctness. Furthermore, we develop a multi-stage evaluation pipeline that systematically assesses model-generated specifications by combining theorem proving and comprehensive testing, providing a practical and robust way to score their soundness and completeness against our ground truth specifications.
Table 1: A comparison of Verina with related prior works on LLMs for code generation and verification. We characterize whether each work supports the three foundational tasks for end-to-end verifiable code generation: CodeGen, SpecGen, ProofGen (Section 4.1). ● means fully supported, ◐ means partially supported, ○ means unsupported. If ProofGen is supported, we specify the proving style: automated theorem proving (ATP) or interactive theorem proving (ITP). For works supporting multiple tasks, we annotate if these tasks are supported in a modular and composable manner. Overall, Verina offers more comprehensive and high-quality benchmarking compared to prior works.
CodeGen SpecGen ProofGen Proving Style Compositionality Language
Benchmarks HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021) Python
Dafny-Synthesis (Misu et al., 2024) ATP Dafny
DafnyBench (Loughridge et al., 2025) ATP Dafny
miniCodeProps (Lohn & Welleck, 2024) ITP Lean
FVAPPS (Dougherty & Mehta, 2025) ITP Lean
Techniques nl2postcond (Endres et al., 2024) Python, Java
Clover (Sun et al., 2024) ATP Dafny
AlphaVerus (Aggarwal et al., 2024) ATP Rust
AutoSpec (Wen et al., 2024) ATP C/C++
SpecGen (Ma et al., 2025) ATP Java
SAFE (Chen et al., 2024) ATP Rust
AutoVerus (Yang et al., 2025) ATP Rust
Laurel (Mugnier et al., 2025) ATP Dafny
Pei et al. (2023) ATP Java
Baldur (First et al., 2023), Selene (Zhang et al., 2024) ITP Isabelle
Rango (Thompson et al., 2025), PALM (Lu et al., 2024) ITP Coq
Verina ITP Lean
@@
The high-quality samples and robust metrics of Verina establish it as a rigorous platform for evaluating verifiable code generation. On Verina, we conduct a thorough experimental evaluation of eight state-of-the-art general-purpose LLMs and three LLMs or agentic frameworks specialized in theorem proving. Our results reveal that even the top-performing general-purpose LLM, OpenAI o4-mini (OpenAI, ), struggles with verifiable code generation, producing only 61.4% correct code solutions, 51.0% sound and complete specifications, and 3.6% successful proof in one trial. Among theorem-proving LLMs, the best model, Goedel Prover V2 32B (Lin et al., 2025), achieved an 11.2% proof success rate in one trial. Interestingly, iterative refinement using Lean compiler feedback can increase the proof success rate to 20.1% with 64 refinement steps. However, this approach significantly raises costs and the success rate remains low. These findings underscore the challenges of verifiable code generation and highlight the critical role of Verina in advancing the field.
@@ 2 Background and Related Work
Task support for verifiable code generation. Writing code, specifications, and proofs for a verified software component is time-consuming when done manually. Although various studies have explored using LLMs to automate these tasks, they primarily focus on individual aspects, failing to capture the full spectrum of verifiable code generation. Benchmarks like HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) have sparked impressive progress on LLM-based code generation but do not handle formal specifications or proofs. Many verification-focused efforts target only one or two tasks, while assuming the other elements are provided by the human user. For example, DafnyBench (Loughridge et al., 2025) and miniCodeProps (Lohn & Welleck, 2024) are two benchmarks designed exclusively for proof generation. Moreover, AutoSpec (Wen et al., 2024) and SpecGen (Ma et al., 2025) infer specifications and proofs from human-written code.
@@
To the best of our knowledge, Dafny-Synthesis (Misu et al., 2024) and Clover (Sun et al., 2024) are the only two works that cover all three tasks, like Verina. However, they target automated theorem proving using Dafny (Leino, 2010), while Verina leverages interactive theorem proving in Lean. Moreover, they have relatively small numbers of human-written samples (50 and 62 respectively). In contrast, Verina provides 189 high-quality samples that are manually validated and undergo rigorous quality assurance (Section 3.2).
@@
Automated and interactive theorem proving. A major challenge in formal verification and verifiable code generation lies in tooling. Verification-oriented languages like Dafny (Leino, 2010) and Verus (Lattuada et al., 2023) leverage SMT solvers for automated theorem proving (De Moura & Bjørner, 2008; Barrett & Tinelli, 2018) and consume only proof hints, such as loop invariants (Pei et al., 2023) and assertions (Mugnier et al., 2025). However, SMT solvers handle only limited proof domains and behave as black boxes, which can make proofs brittle and hard to debug (Zhou et al., 2023). Interactive theorem proving (ITP) systems like Lean provide a promising target for verifiable code generation with LLMs. ITPs support constructing proofs with explicit intermediate steps. This visibility enables LLMs to diagnose errors, learn from unsuccessful steps, and iteratively refine their proofs.Recent work shows that LLMs can generate proofs at human level in math competitions (Google DeepMind, 2024). Prior verification benchmarks in Lean include miniCodeProps (Lohn & Welleck, 2024) and FVAPPS (Dougherty & Mehta, 2025). miniCodeProps translates 201 Haskell programs and their specifications into Lean but is designed for proof generation only. FVAPPS contains 4,715 Lean programs with LLM-generated specifications from a fully automated pipeline that lacks human validation and quality control. In contrast, Verina provides human-verified samples and captures all three foundational tasks in verifiable code generation.
@@
Task compositionality. A key strength of Verina is its modular design, which enables flexible evaluation of not only individual tasks but also their combinations (Section 4.2). This compositionality captures diverse real-world scenarios—from specification-guided code generation to end-to-end verifiable code generation—enabling a comprehensive assessment of different aspects of verifiable code generation. This modularity also facilitates targeted research on specific weaknesses, such as improving proof generation. On the contrary, all other prior works lack full compositionality. For example, Dafny-Synthesis (Misu et al., 2024) and Clover (Sun et al., 2024) mix specification and proof generation into a single task, lacking support for separate evaluation of each.
@@
5-- Code implementation
6def removeElement (s : Array Int) (k : Nat) (h_precond : removeElement_pre s k) : Array Int :=
@@
8-- Pre-condition
9def removeElement_pre (s : Array Int) (k : Nat) : Prop :=
@@
11-- Post-condition
12def removeElement_post (s : Array Int) (k : Nat) (res: Array Int) (h_precond : removeElement_pre s k)
13 : Prop :=
14 res.size = s.size - 1 -- Only one element is removed
15 (∀ i, i < k res[i]! = s[i]!) -- The elements before index k remain unchanged
16 -- The elements after index k are shifted by one position
17 (∀ i, i < res.size i k res[i]! = s[i + 1]!)
18-- Proof (proof body omitted for brevity)
19theorem removeElement_spec (s: Array Int) (k: Nat) (h_precond : removeElement_pre s k) :
20 removeElement_post s k (removeElement s k h_precond) h_precond := by sorry
21-- Test cases
22(s : #[1, 2, 3, 4, 5]) (k : 2) (res : #[1, 2, 4, 5]) -- Positive test with valid inputs and output
23-- Negative test cases
24(s : #[1, 2, 3, 4, 5]) (k : 5) -- Inputs violate the pre-condition at Line 12
25(s : #[1, 2, 3, 4, 5]) (k : 2) (res : #[1, 2, 4]) -- Output violates the post-condition at Line 16
26(s : #[1, 2, 3, 4, 5]) (k : 2) (res : #[2, 2, 4, 5]) -- Output violates the post-condition at Line 17
27(s : #[1, 2, 3, 4, 5]) (k : 2) (res : #[1, 2, 4, 4]) -- Output violates the post-condition at Line 18
@@ 3.1 Overview and Data Format
  • Code (Line 5–7): ground truth code implementation that solves the programming problem.
  • Specification (Line 8–17): ground truth formal specification for the programming problem. It consists of a pre-condition, which states properties the inputs must satisfy, and a post-condition, which states desired relationship between inputs and outputs.
  • Proof (Optional, Line 18–20): formal proof establishing that the code satisfies the specification. Ground truth proofs are optional in Verina, as they are not required for evaluation. Model-generated proofs can be checked by Lean directly. Nevertheless, we invest significant manual effort in writing proofs for 46 out of 189 examples as they help quality assurance (Section 3.2).
  • Test suite (Line 21–27): a comprehensive suite of both positive and negative test cases. Positive tests are valid input-output pairs that meet both the pre-condition and the post-condition. Negative tests are invalid inputs-output pairs, which means either the inputs violate the pre-condition or the output violates the post-condition. These test cases are useful for evaluating model-generated code and specifications, as detailed in Section 4.1. They are formatted in Lean during evaluation.